OpenStudy (anonymous):

5(9 − 3) = w • 9 – 5 • 3 w=

OpenStudy (texaschic101):

5(9-3) = w * 9 - 5 * 3 --- do the parenthesis first 5(6) = w * 9 - 5 * 3 -- now multiply 30 = 9w - 15 --- add 15 to both sides 30 + 15 = 9w --simplify 45 = 9w -- now divide both sides by 9 5 = w check... 5(9-3) = 5 * 9 - 5 * 3 5(6) = 45 - 15 30 = 30 (correct) so w = 5 any questions ?

OpenStudy (anonymous):

Whoah that was a big equation!